# TerraMind: Large-Scale Generative Multimodality for Earth Observation

### v3

@@ 10.3 Finetuning comparisons with baseline models
Since the first approaches to foundation models for Earth observations, experts in the field discuss on the usability of such models compared to task-specific models that are trained for each application individually. Recent benchmark results suggested that task-specific models, like U-Nets, often outperform finetuned GFMs [49]. We therefore additionally investigate how TerraMind compares with task-specific U-Nets and ViT models following the PANGAEA evaluation protocol in Table 6. As adviced by the authors of PANGAEA, we again report results on nine of the eleven datasets as we could not reproduce the performance on the remaining two datasets. The task-specific models are trained from scratch for each individual task, while all GFMs including TerraMind are finetuned with a frozen encoder and an UperNet head. For reasons of consistency, we report TerraMind results based on full finetuning instead of TiM tuning. Yet, for several tasks, we expect slightly higher results when applying TiM tuning for TerraMind. Overall, our results demonstrate that TerraMindv1-B outperforms task-specific UNet and ViT models across the PANGAEA benchmark in both unimodal and multimodal settings by 1pp avg. mIoU and 4pp avg. mIoU respectively. In multimodal settings, the improvement peaks to 4.5pp improvement of TerraMindv1-B over task-specific U-Nets. To the best of our knowledge, this is the first time a GFM model outperforms task-specific models on a global benchmark.
@@ 10.4 Comparing generations and pseudo-labels
We evaluate the model generations for modalities where we used pseudo-labels as input data. For example, in initial experiments with TerraMind.v0.1, we leverage Google’s DynamicWorld model to pseudo-label LULC maps which we use as input to the model. In the following experiment in Table 12, we test the performance of the DynamicWorld model against the generations of TerraMind. Overall, we observe that while finetuned TerraMind.v0.1-B outperforms DynamicWorld, the generation of TerraMind does not surpass the inference results of DynamicWorld.

### v4

@@ 10.3 Finetuning comparisons with baseline models
Since the first approaches to foundation models for Earth observations, experts in the field discuss on the usability of such models compared to task-specific models that are trained for each application individually. Recent benchmark results suggested that task-specific models, like U-Nets, often outperform finetuned GFMs [49]. We therefore additionally investigate how TerraMind compares with task-specific U-Nets and ViT models following the PANGAEA evaluation protocol in Table 6. As adviced by the authors of PANGAEA, we again report results on nine of the eleven datasets as we could not reproduce the performance on the remaining two datasets. The task-specific models are trained from scratch for each individual task, while all GFMs including TerraMind are finetuned with a frozen encoder and an UperNet head. Overall, our results demonstrate that TerraMindv1-B outperforms task-specific UNet and ViT models across the PANGAEA benchmark in both unimodal and multimodal settings by 1pp avg. mIoU and 4pp avg. mIoU respectively. In multimodal settings, the improvement peaks to 4.5pp improvement of TerraMindv1-B over task-specific U-Nets. To the best of our knowledge, this is the first time a GFM model outperforms task-specific models on a global benchmark.
@@ 10.4 Comparing generations and pseudo-labels
We evaluate the model generations for modalities where we used pseudo-labels as input data. For example, in initial experiments with TerraMindv1-B-single, we leverage Google’s DynamicWorld model to pseudo-label LULC maps which we use as input to the model. In the following experiment in Table 12, we test the performance of the DynamicWorld model against the generations of TerraMind. Overall, we observe that while finetuned TerraMindv1-B-single outperforms DynamicWorld, the generation of TerraMind does not surpass the inference results of DynamicWorld.